# algorithmic

### v2

@@ 5.1 Implementation
The backbone network architecture for our mixture model is a multilayer perceptron (MLP) –- a fully connected feedforward network with several hidden layers and a nonlinear activation function for all nodes. The use of an MLP is consistent with the literature, showing strong predictive performance (Katzman et al., 2018; Nagpal et al., 2021; Lillelund et al., 2024b). For all experiments, we use a single layer with {32, 64, 128} nodes and the ReLU6 activation function. We train MENSA with the Adam optimizer (Kingma and Ba, 2017), and use learning rates of $\lambda\in\{1e-3,5e-4,1e-4\}$ for the MLP, and $\Psi\in\{1,3,5\}$ number of Weibull distributions for each event. We add regularization using weight decay of {1e-3, 5e-3, 1e-2} and dropout rates of {0, 0.1, 0.25, 0.5} with batch normalization. Training is done using minibatches, i.e., stochastic subsets of the training dataset with sizes of {32, 64, 128}. We use hyperparameter optimization on the validation set, for each of these hyperparameters We provide additional details on implementation in Appendix F.
@@
Concerning the empirical experiments, we evaluate all models over 10 experiments by applying different survival baselines to six survival datasets with different random seed (0-9) to ensure variability. In each experiment, we follow Sechidis et al. (2011) and first split the dataset into training (70%), validation (10%), and test (20%) sets in a stratified manner wrt the event indicators and event times across the $K$ events given a random seed. The stratification ensures a balanced representation of the time-to-event information and outcome across the data splits. After splitting, we impute missing values using the sample mean for real-valued features or the mode for categorical features based on the training set only. We then apply a $z$-score data normalization to speed up the training process and improve stability. Lastly, we encode categorical features using a one-hot encoding strategy.
@@ 5.2 Datasets
This section briefly describes the survival datasets in this work. Table 2 shows an overview of the datasets. We provide further details on the data and preprocessing in Appendix D.
@@
Synthetic: Following Foomani et al. (2023), we create synthetic, single-event datasets from a data generation process (DGP) with a linear or nonlinear activation function. To evaluate our model under dependent censoring, i.e., where the event and censoring times are not independent, we sample $T_{E}$ and $T_{C}$ R.V.s from a so-called copula. A copula $C_{\theta}(u,v)$ links R.V.s by specifying their dependence structure (Emura and Chen, 2018, Ch. 1). Applied here, it captures the dependency (Kendall’s $\tau$) between $T_{E}$ and $T_{C}$ by sampling the marginals from a (survival) copula. If $\tau=0$, the marginals are generated as independent uniform R.V.s, otherwise, they are generated using an Archimedean copula. We provide more details about this data generation process in Appendix D.
@@
SEER: The U.S. Surveillance, Epidemiology, and End Results (SEER) dataset (Ries et al., 2003) is a comprehensive collection of cancer patient data from approximately 49% of the U.S. population. It provides information about survival times for cancer patients post diagnosis. We select a cohort of 19,246 newly-diagnosed patients (first year of disease) with 17 features, which include demographics and tumor characteristics. For the single-event case, we consider the outcome of death due to breast cancer, and for the competing-risks case, we consider the outcomes of death due to breast cancer and death due to heart failure – and consider other outcomes as censored.
@@
MIMIC-IV: The Medical The Information Mart for Intensive Care (MIMIC-IV) dataset by Johnson et al. (2023) comprises electronic health records from patients admitted to intensive care units at the Beth Israel Deaconess Medical Center in the U.S., covering the years from 2008 to 2019. Using the method described in Gupta et al. (2022), we extract 1,672 static features from the MIMIC-IV dataset, for 26,236 patients whose age is between 60 and 65 years. To refine our feature set, we apply the unicox feature selection method (Qi et al., 2022; Simon et al., 2003), reducing the number of features to the 100 that appear to be the most relevant (see Appendix D for details). For the single-event case, we consider death since hospital admission and for the multi-event case, we consider the outcomes of acute respiratory failure (ARF), shock and death.
@@
Rotterdam: The Rotterdam dataset (Royston and Altman, 2013) includes 2,982 primary breast cancer patients from the Rotterdam tumor bank, with 1,546 having node-positive disease. The follow-up time ranged from 1 to 231 months (median 107 months). The dataset has 10 features, which include demographics, tumor characteristics, and treatment information. In the competing-risks case, we consider the outcome of recurrence-free survival time (RFS), which is defined as the time from primary surgery to the earlier of disease recurrence or death from any cause. In the multi-event case, we consider the time to death or disease recurrence, where death censors disease recurrence, but disease recurrence does not censor death. Figure 7 shows the event trajectories in the Rotterdam dataset.
@@
PRO-ACT: The Pooled Resources Open-Access Clinical Trials (PRO-ACT) [^1] Data used in the preparation of this article were obtained from the Pooled Resource Open-Access ALS Clinical Trials (PRO-ACT) Database. The data available in the PRO-ACT Database have been volunteered by PRO-ACT Consortium members. database by Atassi et al. (2014) is the largest ALS dataset in the world, which includes patient demographics, lab and medical records and family history of over 11,600 ALS patients, where ALS disease state is evaluated by the ALS functional rating scale revised (ALSFRS-R). The assessment contains 12 questions and a maximum score of 48, each representing physical functionality, e.g., the ability to walk. Each ALSFRS-R score is between 0 and 4, where 0 represents complete inability with regard to the function, and 4 represents normal function. We use the ALSFRS-R scores to formulate four events: “Speech”, “Swallowing”, “Handwriting” and “Walking”. We consider an event to have occurred if a patient scores 2 or lower on the ALSFRS-R scale during any subsequent follow-up visitation following their baseline visitation. The events can occur in any order and are not mutually exclusive. ALSFRS-R scores are generally monotonic decreasing, but due to the subjectivity of the assessment, scores can occasionally increase post-baseline. We use a maximum follow-up time of 500 days and exclude patients with no recorded history.
EBMT: The European Society for Blood and Marrow Transplantation (EBMT) dataset (Marta Fiocco, 2008) contains records of 2,279 patients who underwent blood and marrow transplantation between 1985 and 1998 for acute lymphoid leukemia. The dataset has several features, including gender mismatch between donor and recipient, age at transplant, whether any GvHD prophylaxis was administered, and year of transplantation. We consider five events: platelet recovery from acute-graft-versus-host disease (AGvHD), adverse effects after transplantation, recovery and adverse effects, disease relapse and death in remission. If AGvHD occurs, it will occur shortly after the transplant, usually within the first 100 days. Figure 8 shows the event trajectories in the EBMT dataset.
@@ 5.3 Baseline models
CoxPH: The Cox Proportional Hazards model (CoxPH) is a shallow semiparametric model for survival analysis (Cox, 1972). The conditional individual hazard function assumed by the CoxPH model is $h\left(t|\bm{x}^{(i)}\right)=h_{0}\left(t\right)\exp\left(f\left(\bm{\bm{\theta}},\bm{x}^{(i)}\right)\right)$, where $h_{0}\left(t\right)$ is the baseline hazard at $t$, $f\left(\bm{\bm{\theta}},\bm{x}^{(i)}\right)$ is a risk score, and $f$ is a linear function. Its standard MLE of $\hat{\bm{\bm{\theta}}}$ optimizes the so-called Cox partial log-likelihood (Cox, 1972).
@@
GBSA: GBSA (Ridgeway, 1999) is a shallow likelihood-based boosting model that optimizes an $\ell_{2}$-norm penalized Cox partial log-likelihood (Cox, 1972). It updates candidate variables flexibly at each boosting step using an offset-based gradient boosting method, unlike traditional gradient boosting, which updates either a single component or all features for fitting the gradient.
@@
RSF: Random Survival Forests (RSF) (Ishwaran et al., 2008) extends decision trees for survival analysis and recursively splits training data based on some survival-specific criterion, e.g., log-rank splitting, with the goal of maximizing separation between survival times across nodes.
@@
DeepSurv: DeepSurv (Katzman et al., 2018) is a deep continuous-time survival model of the form $h\left(t|\bm{x}^{(i)}\right)=h_{0}\left(t\right)\exp\left(f\left(\bm{\theta},\bm{x}^{(i)}\right)\right)$, where $f\left(\bm{\theta},\bm{x}^{(i)}\right)$ denotes a risk score as a nonlinear function of the features, i.e.$f\left(\bm{\theta},\bm{x}^{(i)}\right)=\sigma(\bm{x}^{(i)}\bm{\theta})$. The MLE for $\hat{\bm{\theta}}$ is derived by numerically maximizing the partial Cox log-likelihood.
@@
DeepHit: DeepHit (Lee et al., 2018) is a deep discrete-time model designed to handle competing risks, where the occurrence of one event prevents the occurrence of another, as it assumes $\sum_{t=1}^{T}\sum_{k=1}^{K}p_{k}[t]=1$. It features a shared subnetwork among all events and cause-specific subnetworks for each competing event. The model outputs a probability distribution for each discrete time bin and each competing event via a softmax layer. This architecture allows DeepHit to estimate the joint distribution of survival times and events directly without making assumptions about the underlying stochastic process. It accounts for dependencies among events in $\bm{\theta}$ but not for individuals who are event-free at time $T$.
@@
Hierarchical: The hierarchical model (Tjandra et al., 2021) is a deep discrete-time model designed to handle multi-event. The model learns the individual survival distribution by iteratively predicting the probability of an event at increasingly finer time scales. During training, the model uses a composite likelihood function using ranking penalties that provide supervision across the entire event horizon.
@@
MTLR: The Multi-Task Logistic Regression (MTLR) model (Yu et al., 2011) is a discrete-time survival model that predicts the survival distribution using a sequence of dependent logistic regressors. The MTLR model has been extended to the competing-risks case as the Deep-CR MTLR model by Kim et al. (2021).
@@
DSM: Deep Survival Machines (DSM) (Nagpal et al., 2021) is a deep continuous-time model that estimates the survival distribution as a mixture of some fixed $k$ parametric distributions. The parameters of these mixture distributions and the mixing weights are estimated using an MLP.
@@ 5.4 Evaluation metrics
For the single-event datasets, we evaluate discrimination using the traditional Harrell’s concordance index (CI) (Harrell Jr. et al., 1982). In the competing-risks and multi-event datasets, we report the Global CI as the average of the CIs across events, and the Local CI, as the average of the CIs for each individual. Thus, these two discrimination metrics evaluate global consistency, i.e., how well does the model discriminate individuals within an event, and local consistency, i.e., how well does the model discriminate events within an individual. Figure 9 shows an example of this with respect to the survival curve. Visually, global consistency is satisfied between instances if the instance-specific curves align with the observed events. Local consistency is satisfied within an instance if the event-specific curves align with the observed events. We evaluate the accuracy of the predicted time-to-event estimates using mean absolute error (MAE) with a margin loss for censored individuals (mMAE(Haider et al., 2020) and the integrated Brier Score (IBS(Graf et al., 1999). To evaluate calibration, we report the Distribution-Calibration (D-Cal) score (Haider et al., 2020), which measures how well the survival function is calibrated for each event. Lastly, we evaluate the bias induced by dependent censoring using the Survival-$\ell_{1}$ error (Foomani et al., 2023), which is the $\ell_{1}$ distance between the ground truth survival curve and the estimated survival curve. See Appendix B for more details about the evaluation metrics.
@@ 6.1 Single-event prediction
We first consider the single-event case ($K=1$), and evaluate our method under dependent censoring. Dependent censoring occurs when the likelihood of being censored is related to the event time, which violates the key assumption of independent censoring behind covariate-based survival models (Emura and Chen, 2018, Ch. 3). Models that assume independence may therefore underestimate or overestimate the survival function due to the mismatch between the observed and true survival distributions. To investigate our method under dependent censoring, we train MENSA and baseline models on synthetic datasets with either linear or nonlinear Weibull hazards and with various degrees of dependence from either a Frank or Clayton copula. Figure 10 shows the performance results in terms of the bias (Survival-$\ell_{1}$) for MENSA and baseline models under dependent censoring. In the linear case, we see that MENSA has one of the lowest Survival-$\ell_{1}$ errors, and that its error is stable as a function of the dependence. In the nonlinear case, MENSA has above average error using a Frank copula but the second lowest error using a Clayton copula. This suggests that MENSA is robust to dependent censoring in the linear case under the assumption of dependence from a Frank or Clayton copula, and has lower estimation bias in the nonlinear case than all rival methods, except DeepHit, under a Clayton copula.
@@
Table 3: Single-event prediction results. We report the mean ($\pm$ SD.) averaged over 10 experiments. D-calibration counts the number of times the respective model was D-calibrated. To improve readability, we have multiplied the CI and IBS by 100 and divided the mMAE results in MIMIC-IV by 100.
@@ 6.2 Competing-risks prediction
We now consider the competing-risks case ($K>1$), where the $K$ events are mutually exclusive –- that is, each instance can have (at most) one event. We evaluate MENSA in the SEER and Rotterdam test sets. Table 4 shows the prediction results. In the SEER dataset, MENSA achieves the best (highest) global concordance index and the third-best (third-highest) local concordance index. MENSA also shines in giving accurate time-to-event predictions in this dataset, as MENSA has the best (lowest) integrated Brier Score and best (lowest) margin MAE, and achieves perfect D-calibration in all 20 experiments. Again, this highlights MENSA’s applicability in multiple event scenarios, where our interest extends beyond ranking and predicting individuals at risk for a single event to include multiple events. In the Rotterdam dataset, MENSA is again very competitive, as it achieves the best (highest) global concordance index and the second-best (second-highest) local concordance index, slightly behind DeepHit in the latter case. Although MENSA’s integrated Brier Score of 15.03 is slightly behind several competitors, it gives D-calibrated survival curves in 18/20 experiments and a margin MAE of 49.56, which is better than all rival models, except DeepHit. As evidenced by the calibration results, MENSA more often give calibrated survival curves than literature baselines in these two datasets.
@@
Table 4: Competing risks prediction results. We report the mean ($\pm$ SD.) averaged over 10 experiments. D-calibration reports the number of experiments in which the respective model was D-calibrated. To improve readability, we have multiplied the CI and IBS results by 100.

### v3

@@ 5.1 Implementation
The backbone network architecture for our mixture model is a multilayer perceptron (MLP) –- a fully connected feedforward network with several hidden layers and a nonlinear activation function for all nodes. The use of an MLP is consistent with the literature, showing strong predictive performance [23, 36, 33]. For all experiments, we use a single layer with {32, 64, 128} nodes and the ReLU6 activation function. We train MENSA with the Adam optimizer [25], and use learning rates of $\lambda\in\{1e-3,5e-4,1e-4\}$ for the MLP, and $\Psi\in\{1,3,5\}$ number of Weibull distributions for each event. We add regularization using weight decay of {1e-3, 5e-3, 1e-2} and dropout rates of {0, 0.1, 0.25, 0.5} with batch normalization. Training is done using minibatches, i.e., stochastic subsets of the training dataset with sizes of {32, 64, 128}. We use hyperparameter optimization on the validation set, for each of these hyperparameters We provide additional details on implementation in Appendix F.
@@
Concerning the empirical experiments, we evaluate all models over 10 experiments by applying different survival baselines to six survival datasets with different random seed (0-9) to ensure variability. In each experiment, we follow Sechidis et al. [46] and first split the dataset into training (70%), validation (10%), and test (20%) sets in a stratified manner wrt the event indicators and event times across the $K$ events given a random seed. The stratification ensures a balanced representation of the time-to-event information and outcome across the data splits. After splitting, we impute missing values using the sample mean for real-valued features or the mode for categorical features based on the training set only. We then apply a $z$-score data normalization to speed up the training process and improve stability. Lastly, we encode categorical features using a one-hot encoding strategy.
@@ 5.2 Datasets
This section briefly describes the survival datasets in this work. Table 2 shows an overview of the datasets. We provide further details about the data in Appendix D.
@@
Synthetic: Following Foomani et al. [12], we create synthetic, single-event datasets from a data generation process (DGP) with a linear or nonlinear activation function. To evaluate our model under dependent censoring, i.e., where the event and censoring times are not independent, we sample $T_{E}$ and $T_{C}$ R.V.s from a so-called copula. A copula $C_{\theta}(u,v)$ links R.V.s by specifying their dependence structure [8, Ch. 1]. Applied here, it captures the dependency (Kendall’s $\tau$) between $T_{E}$ and $T_{C}$ by sampling the marginals from a (survival) copula. If $\tau=0$, the marginals are generated as independent uniform R.V.s, otherwise, they are generated using an Archimedean copula. We provide more details about this data generation process in Appendix D.
@@
SEER: The U.S. Surveillance, Epidemiology, and End Results (SEER) dataset [43] is a comprehensive collection of cancer patient data from approximately 49% of the U.S. population. It provides information about survival times for cancer patients post diagnosis. We select a cohort of 19,246 newly-diagnosed patients (first year of disease) with 17 features, which include demographics and tumor characteristics. For the single-event case, we consider the outcome of death due to breast cancer, and for the competing-risks case, we consider the outcomes of death due to breast cancer and death due to heart failure – and consider other outcomes as censored.
@@
MIMIC-IV: The Medical The Information Mart for Intensive Care (MIMIC-IV) dataset by Johnson et al. [21] comprises electronic health records from patients admitted to intensive care units at the Beth Israel Deaconess Medical Center in the U.S., covering the years from 2008 to 2019. Using the method described in Gupta et al. [15], we extract 1,672 static features from the MIMIC-IV dataset, for 26,236 patients whose age is between 60 and 65 years. To refine our feature set, we apply the unicox feature selection method [39, 47], reducing the number of features to the 100 that appear to be the most relevant (see Appendix D for details). For the single-event case, we consider death since hospital admission and for the multi-event case, we consider the outcomes of acute respiratory failure (ARF), shock and death.
@@
Rotterdam: The Rotterdam dataset [45] includes 2,982 primary breast cancer patients from the Rotterdam tumor bank, with 1,546 having node-positive disease. The follow-up time ranged from 1 to 231 months (median 107 months). The dataset has 10 features, which include demographics, tumor characteristics, and treatment information. In the competing-risks case, we consider the outcome of recurrence-free survival time (RFS), which is defined as the time from primary surgery to the earlier of disease recurrence or death from any cause. In the multi-event case, we consider the time to death or disease recurrence, where death censors disease recurrence, but disease recurrence does not censor death. Figure 9 shows the event trajectories in the Rotterdam dataset.
@@
PRO-ACT: The Pooled Resources Open-Access Clinical Trials (PRO-ACT) [^1] Data used in the preparation of this article were obtained from the Pooled Resource Open-Access ALS Clinical Trials (PRO-ACT) Database. The data available in the PRO-ACT Database have been volunteered by PRO-ACT Consortium members. database by Atassi et al. [4] is the largest ALS dataset in the world, which includes patient demographics, lab and medical records and family history of over 11,600 ALS patients, where ALS disease state is evaluated by the ALS functional rating scale revised (ALSFRS-R). The assessment contains 12 questions and a maximum score of 48, each representing physical functionality, e.g., the ability to walk. Each ALSFRS-R score is between 0 and 4, where 0 represents complete inability with regard to the function, and 4 represents normal function. We use the ALSFRS-R scores to formulate four events: “Speech”, “Swallowing”, “Handwriting” and “Walking”. We consider an event to have occurred if a patient scores 2 or lower on the ALSFRS-R scale during any subsequent follow-up visitation following their baseline visitation. The events can occur in any order and are not mutually exclusive. ALSFRS-R scores are generally monotonic decreasing, but due to the subjectivity of the assessment, scores can occasionally increase post-baseline. We use a maximum follow-up time of 500 days and exclude patients with no recorded history.
EBMT: The European Society for Blood and Marrow Transplantation (EBMT) dataset [35] contains records of 2,279 patients who underwent blood and marrow transplantation between 1985 and 1998 for acute lymphoid leukemia. The dataset has several features, including gender mismatch between donor and recipient, age at transplant, whether any GvHD prophylaxis was administered, and year of transplantation. We consider five events: platelet recovery from acute-graft-versus-host disease (AGvHD), adverse effects after transplantation, recovery and adverse effects, disease relapse and death in remission. If AGvHD occurs, it will occur shortly after the transplant, usually within the first 100 days. Figure 10 shows the event trajectories in the EBMT dataset.
@@ 5.3 Baseline models
CoxPH: The Cox Proportional Hazards model (CoxPH) is a shallow semiparametric model for survival analysis [6]. The conditional individual hazard function assumed by the CoxPH model is $h\left(t|\bm{x}^{(i)}\right)=h_{0}\left(t\right)\exp\left(f\left(\bm{\bm{\theta}},\bm{x}^{(i)}\right)\right)$, where $h_{0}\left(t\right)$ is the baseline hazard at $t$, $f\left(\bm{\bm{\theta}},\bm{x}^{(i)}\right)$ is a risk score, and $f$ is a linear function. Its standard MLE of $\hat{\bm{\bm{\theta}}}$ optimizes the so-called Cox partial log-likelihood.
@@
GBSA: GBSA [42] is a shallow likelihood-based boosting model that optimizes an $\ell_{2}$-norm penalized Cox partial log-likelihood. It updates candidate variables flexibly at each boosting step using an offset-based gradient boosting method, unlike traditional gradient boosting, which updates either a single component or all features for fitting the gradient.
@@
RSF: Random Survival Forests (RSF) [19] extends decision trees for survival analysis and recursively splits training data based on some survival-specific criterion, e.g., log-rank splitting, with the goal of maximizing separation between survival times across nodes.
@@
DeepSurv: DeepSurv [23] is a deep continuous-time survival model of the form $h\left(t|\bm{x}^{(i)}\right)=h_{0}\left(t\right)\exp\left(f\left(\bm{\theta},\bm{x}^{(i)}\right)\right)$, where $f\left(\bm{\theta},\bm{x}^{(i)}\right)$ denotes a risk score as a nonlinear function of the features, i.e.$f\left(\bm{\theta},\bm{x}^{(i)}\right)=\sigma(\bm{x}^{(i)}\bm{\theta})$. The MLE for $\hat{\bm{\theta}}$ is derived by numerically maximizing the partial Cox log-likelihood.
@@
DeepHit: DeepHit [30] is a deep discrete-time model designed to handle competing risks, where the occurrence of one event prevents the occurrence of another, as it assumes $\sum_{t=1}^{T}\sum_{k=1}^{K}p_{k}[t]=1$. It features a shared subnetwork among all events and cause-specific subnetworks for each competing event. The model outputs a probability distribution for each discrete time bin and each competing event via a softmax layer. This architecture allows DeepHit to estimate the joint distribution of survival times and events directly without making assumptions about the underlying stochastic process. It accounts for dependencies among events in $\bm{\theta}$ but not for individuals who are event-free at time $T$.
@@
Hierarchical: The hierarchical model [51] is a deep discrete-time model designed to handle multiple event. The model learns the individual survival distribution by iteratively predicting the probability of an event at increasingly finer time scales. During training, the model uses a composite likelihood function using ranking penalties that provide supervision across the entire event horizon.
@@
MTLR: The Multi-Task Logistic Regression (MTLR) model [55] is a discrete-time survival model that predicts the survival distribution using a sequence of dependent logistic regressors. The MTLR model has been extended to the competing-risks case as the Deep-CR MTLR model by Kim et al. [24].
@@
DSM: Deep Survival Machines (DSM) [36] is a deep continuous-time model that estimates the survival distribution as a mixture of some fixed $k$ parametric distributions. The parameters of these mixture distributions and the mixing weights are estimated using an MLP.
@@ 5.4 Evaluation metrics
For the single-event datasets, we evaluate discrimination using the traditional Harrell’s concordance index (CI) [17]. In the competing-risks and multi-event datasets, we report the Global CI as the average of the CIs across events, and the Local CI, as the average of the CIs across individuals. Thus, these two discrimination metrics evaluate global consistency, i.e., how well does the model discriminate individuals within an event, and local consistency, i.e., how well does the model discriminate events within an individual, respectively. Figure 11 shows an example of this with respect to the survival curve. Visually, global consistency is satisfied between instances if the instance-specific curves align with the observed events. Local consistency is satisfied within an instance if the event-specific curves align with the observed events. We evaluate the accuracy of the predicted time-to-event estimates using mean absolute error (MAE) with a margin loss for censored individuals (mMAE[16] and the integrated Brier Score (IBS[14]. To evaluate calibration, we report the Distribution-Calibration (D-Cal) score [16], which measures how well the survival function is calibrated for each event. Lastly for the synthetic experiments, we evaluate the bias induced by dependent censoring using the Survival-$\ell_{1}$ error [12], which is the $\ell_{1}$ distance between the ground truth survival curve and the estimated survival curve. See Appendix B for more details about the evaluation metrics.
@@ 6.1 Single-event prediction
We first consider the single-event case ($K=1$), and evaluate our method under dependent censoring. Dependent censoring occurs when the likelihood of being censored is related to the event time, which violates the key assumption of independent censoring behind covariate-based survival models [8, Ch. 3]. Models that assume independence may therefore underestimate or overestimate the survival function due to the mismatch between the observed and true survival distributions. To investigate our method under dependent censoring, we train MENSA and baseline models on synthetic datasets with either linear or nonlinear Weibull hazards and with various degrees of dependence from either a Frank or Clayton copula. Figure 12 shows the performance results in terms of the bias (Survival-$\ell_{1}$) for MENSA and baseline models under dependent censoring. In the linear case, we see that MENSA has one of the lowest Survival-$\ell_{1}$ errors, and that its error is stable as a function of the dependence. In the nonlinear case, MENSA has above average error using a Frank copula but the second lowest error using a Clayton copula. This suggests that MENSA is robust to dependent censoring in the linear case under the assumption of dependence from a Frank or Clayton copula, and has lower estimation bias in the nonlinear case than all rival methods, except DeepHit, under a Clayton copula.
@@
Table 3: Single-event prediction results. We report the mean ($\pm$ SD.) averaged over 10 experiments. D-calibration counts the number of times the respective model was D-calibrated. To improve readability, CI and IBS were multiplied by 100, and mMAE in MIMIC-IV was divided by 100.
@@ 6.2 Competing-risks prediction
We now consider the competing-risks case ($K>1$), where the $K$ events are mutually exclusive –- that is, each instance can have (at most) one event. We evaluate MENSA in the SEER and Rotterdam test sets. Table 4 shows the prediction results. In the SEER dataset, MENSA achieves the best (highest) global concordance index and the third-best (third-highest) local concordance index. MENSA also shines in giving accurate time-to-event predictions in this dataset, as MENSA has the best (lowest) integrated Brier Score and best (lowest) margin MAE, and achieves perfect D-calibration in all 20 experiments. Again, this highlights MENSA’s applicability in multi-event scenarios, where our interest extends beyond ranking and predicting individuals at risk for a single event to include multiple events. In the Rotterdam dataset, MENSA is again very competitive, as it achieves the best (highest) global concordance index and the second-best (second-highest) local concordance index, slightly behind DeepHit in the latter case. Although MENSA’s integrated Brier Score of 15.03 is slightly behind several competitors, it gives D-calibrated survival curves in 18/20 experiments and a margin MAE of 49.56, which is better than all rival models, except DeepHit. As evidenced by the calibration results, MENSA more often give calibrated survival curves than literature baselines in these two datasets.
@@
Table 4: Competing risks prediction results. We report the mean ($\pm$ SD.) averaged over 10 experiments. D-calibration reports the number of experiments in which the respective model was D-calibrated. To improve readability, CI and IBS were multiplied by 100.